OpenStudy (anonymous):

Help me solve these problems of simultaneous equations please :).. 2. X-Y=73 X-2Y=4 3. X+Y=60 3X-2Y=70

OpenStudy (lgbasallote):

x - y = 73 x - 2y = 4 subtract these two equations. what do you get?

OpenStudy (anonymous):

How am i supposed to subtract those two equations??

OpenStudy (anonymous):

YOU want me to subtract the answers or sumn?

OpenStudy (anonymous):

\[x - y = 73\]......(1) \[X-2Y=4\]..........(2) subtract 1 & 2 \[y=69\] \[x=143\]